# HLV-1K: A Large-scale Hour-Long Video Benchmark for Time-Specific Long Video Understanding

### v1

@@ Abstract
Multimodal large language models have become a popular topic in deep visual understanding due to many promising real-world applications. However, hour-long video understanding, spanning over one hour and containing tens of thousands of visual frames, remains under-explored because of 1) challenging long-term video analyses, 2) inefficient large-model approaches, and 3) lack of large-scale benchmark datasets. Among them, in this paper, we focus on building a large-scale hour-long long video benchmark, HLV-1K, designed to evaluate long video understanding models. HLV-1K comprises 1009 hour-long videos with 14,847 high-quality question answering (QA) and multi-choice question asnwering (MCQA) pairs with time-aware query and diverse annotations, covering frame-level, within-event-level, cross-event-level, and long-term reasoning tasks. We evaluate our benchmark using existing state-of-the-art methods and demonstrate its value for testing deep long video understanding capabilities at different levels and for various tasks. This includes promoting future long video understanding tasks at a granular level, such as deep understanding of long live videos, meeting recordings, and movies.
@@ I Introduction
Figure 1: HLV-1K: (a) Video category distribution, (b) Video duration distribution, and (c) Duration distribution of time-specific query.
@@
  • Diverse tasks across different levels. HLV-1K includes both QA and MCQA tasks, encompassing a variety of reasoning tasks. This is in contrast to other long video benchmarks, such as Video-MME [15] and LVBench [16], which primarily use only MCQA in benchmark designing. Additionally, HLV-1K provides reasoning questions based on visual information at the frame-level, within-event-level, cross-event-level, and long-term-level, offering a comprehensive assessment of a model’s capabilities in long-video understanding.
  • Time-specific queries. HLV-1K annotations include precise time information linking QA queries to video content, ensuring models are evaluated on their ability to reason about temporal relationships and long-term dependencies.
@@ II-A MM-LLMs for Long Video Understanding
Multimodal LLMs integrate LLMs with visual modality encoders and show promising performance in visual understanding tasks, especially image and short video understanding tasks [17, 18, 19, 20]. However, visual understanding for long videos that span minutes or even hours can be challenging due to the varying scenes and visual content, which dramatically increase spatial and temporal details and require capturing long-term correlations over extended time spans [10]. On one hand, most existing long video understanding methods focus on compressing sequential visual frame tokens, providing reasonable performance on minute-level long video understanding tasks [21, 22]. On the other hand, some methods opt to fine-tune models on longer video datasets, such as Moment-10M [23], and employ more efficient visual data compression [24] or long-term information preserving methods [25] to enhance the capability of handling hour-level long video understanding. These approaches aim to balance the need for detailed visual information with the computational efficiency required to process extended video content effectively.
@@ V Conclusion
Appendix 1: Overview of HLV-1K
V-A Annotation Prompts
We extract frame descriptions and further summarize event descriptions using GPT-4o [28]. The prompts used for frame description extraction and event description summarization are shown in Fig. 5.
Figure 5: Prompts for Frame Description and Event Description.
V-B Frame and Event Descriptions
An example of the extracted frame description and summarized event description is shown in Fig. 6. We also provide statistics on the number of frame descriptions and event descriptions for event videos, as well as the word count for event frame descriptions and event descriptions, as shown in Fig. 7.
Figure 6: Examples of Frame Description and Event Description.
Figure 7: Distribution of the number and word of Frame and Event Descriptions.
V-C Source of Multi-level Video Annotations from Frame and Event Descriptions
The multi-level annotations are derived from various sources of visual descriptions, as summarized in Table III.
TABLE III: Source of multi-level video annotations
Level Source
Frame-level Piece of frame description
Within-event-level Piece of event description
Cross-event-level Neighboring event descriptions
Long-term-level All event description within one video
V-D Examples of Multi-task Annotation
Examples of annotations for various tasks, including character understanding, counting, object understanding, spatial relationships, descriptive scenes, speed, object direction, scene understanding, attribute change, time order, temporal relationships, causality, plot understanding, camera direction, and action understanding, are shown in Fig. 8 and Fig. 9.
Figure 8: Examples of Multi-task Annotation (1).
Figure 9: Examples of Multi-task Annotation (2).
VI Appendix 2: Model Evaluation
VI-A Task Prompts
To fairly evaluate the performance of various models, we design two unified prompts for both QA and MCQA tasks, respectively. The prompts used for these two different types of tasks across different models are shown in Fig. 10.
Figure 10: Prompts for $M$-LLMs evaluation on HLV-1K.

### v2

@@ Abstract
Multimodal large language models have become a popular topic in deep visual understanding due to many promising real-world applications. However, hour-long video understanding, spanning over one hour and containing tens of thousands of visual frames, remains under-explored because of 1) challenging long-term video analyses, 2) inefficient large-model approaches, and 3) lack of large-scale benchmark datasets. Among them, in this paper, we focus on building a large-scale hour-long long video benchmark, HLV-1K  [^1] https://github.com/Vincent-ZHQ/HLV-1K, designed to evaluate long video understanding models. HLV-1K comprises 1009 hour-long videos with 14,847 high-quality question answering (QA) and multi-choice question asnwering (MCQA) pairs with time-aware query and diverse annotations, covering frame-level, within-event-level, cross-event-level, and long-term reasoning tasks. We evaluate our benchmark using existing state-of-the-art methods and demonstrate its value for testing deep long video understanding capabilities at different levels and for various tasks. This includes promoting future long video understanding tasks at a granular level, such as deep understanding of long live videos, meeting recordings, and movies.
@@ I Introduction
Figure 1: HLV-1K statistics: (a) Video category distribution, (b) Video duration distribution, and (c) Duration distribution of time-specific queries.
@@
  • Diverse reasoning tasks. HLV-1K includes both QA and MCQA tasks, unlike Video-MME [15] and LVBench [16], which mainly use MCQA. It also introduces reasoning across frame-level, within-event, cross-event, and long-term contexts for comprehensive evaluation.
  • Time-specific queries. Each query in HLV-1K is linked to precise timestamps, enabling robust assessment of models’ temporal reasoning and long-term understanding.
@@ II-A MM-LLMs for Long Video Understanding
Multimodal LLMs integrate LLMs with visual encoders and excel in image and short video understanding [17, 18, 19, 20]. However, long video understanding remains challenging due to increased spatial-temporal complexity and long-term correlations [10]. Existing methods either compress sequential frame tokens for minute-level videos [21, 22] or fine-tune on longer video datasets like Moment-10M [23], leveraging efficient compression [24] or long-term information preservation [25] to enhance hour-level video understanding. These approaches aim to balance visual detail retention with computational efficiency.